OpenStudy (anonymous):

Annabelle's Work (4275)(4372) = 42 + 375 + 2 = 4577 Navene's Work (4275)(4372) = 42⋅375⋅2 = 46710 Are either of them right ?

OpenStudy (anonymous):

No they are not correct. Annabelle's work is incorrect. Navene's work: Also incorrect. \[42 + 375 + 2 =419 = 31,500\]

OpenStudy (anonymous):

\[42\times 375 \times 2 =31,500\]

OpenStudy (anonymous):

Thus, no they're incorrect.